OpenStudy (anonymous):

Really need help! [2.03] Solve: 7x − 2(x + 1) = 6x + 14 −12 12 −16 16 Rachael's brother is 33 years old. He is 3 times Rachael's age 16 years ago. How old is Rachael? 11 27 16 24 2(a + b) = 2a + 2b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

Parth (parthkohli):

The 3rd one: Distributive property. It says: a(x + y) = ax + ay

Parth (parthkohli):

Let's take Rachael's age as x. 33 - 16 = 3(x - 16) Solve the equation.

OpenStudy (anonymous):

got it

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 7x - 2(x + 1) = 6x + 14 }\) \(\Large \color{purple}{\rightarrow 7x - 2x - 2 = 6x + 14 }\) Now, solve.

OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 69. n + (n + 2) = 69 n + (n + 1) = 69 n − (n + 1) = 69 n + 2n = 69 ?

Parth (parthkohli):

Take one integer as n, and then the consecutive integers can be {n, n + 1} or {n, n - 1}

Parth (parthkohli):

We have to find the sum, so add those.

OpenStudy (anonymous):

dont get it add what?

Parth (parthkohli):

n + (n + 1) = 69 n and n + 1 are consecutive integers.

OpenStudy (anonymous):

thanks 1 more left Choose the procedure you would use to solve the equation 4x − 7 = 29 for x. Add seven to both sides, and then multiply both sides by four. Divide both sides by four, and then subtract seven from both sides. Add seven to both sides, and then divide both sides by four. Multiply both sides by four, and then add seven to both sides.